OpenStudy (anonymous):

Convert 8.2 weeks to years!!

OpenStudy (phi):

you multiply by a ratio of \[ \frac{years}{weeks} \] how many weeks in 1 year? there are 52 weeks in 1 year so the ratio is \[ \frac{1 \ year}{52 \ weeks} \] next you write \[ 8.2 \cancel{weeks} \cdot \frac{1\ years}{52\cancel{weeks}} \\= \frac{8.2 \ years}{52} \]

OpenStudy (phi):

you can change that to 82/520 or 41/260 or to a decimal